OpenStudy (anonymous):

What is the limit of (sqrt (x) - 5)/(x-25) as x approaches 25?

hartnn (hartnn):

multiply and divide by (sqrt(x)+5)

OpenStudy (anonymous):

Why is that?

OpenStudy (anonymous):

To rationalize it?

hartnn (hartnn):

use \( (\sqrt x -5)(\sqrt x+5)= x-25\)

hartnn (hartnn):

yup

OpenStudy (anonymous):

\[(x-25)/(x \sqrt{x}+5x-25\sqrt{x}-125\]

hartnn (hartnn):

then u can can out x-25 from numerator and denominator.

OpenStudy (anonymous):

Ah, then you get 1/sqrt{x}+5

OpenStudy (anonymous):

So the limit = 1/10! Wow!

hartnn (hartnn):

?? keep x-25 in denominator as it is.

hartnn (hartnn):

yup , u got it :)

OpenStudy (anonymous):

You're very helpful, thank you!

hartnn (hartnn):

welcome :)